# frame=single, fontsize=, breaklines

### v1

@@ Abstract
The increased use of large language models (LLMs) across a variety of real-world applications calls for automatic tools to check the factual accuracy of their outputs, as LLMs often hallucinate. This is difficult as it requires assessing the factuality of free-form open-domain responses. While there has been a lot of research on this topic, different papers use different evaluation benchmarks and measures, which makes them hard to compare and hampers future progress. To mitigate these issues, we developed OpenFactCheck, a unified framework, with three modules: (iResponseEval, which allows users to easily customize an automatic fact-checking system and to assess the factuality of all claims in an input document using that system, (iiLLMEval, which assesses the overall factuality of an LLM, and (iiiCheckerEval, a module to evaluate automatic fact-checking systems. OpenFactCheck is open-sourced [^1] https://github.com/hasaniqbal777/openfactcheck and publicly released as a Python library [^2] https://pypi.org/project/openfactcheck/ and also as a web service [^3] https://huggingface.co/spaces/hasaniqbal777/OpenFactCheck. A video describing the system is available at https://youtu.be/-i9VKL0HleI.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in generating naturally-sounding answers over a broad range of human inquiries. However, GPT-4o (OpenAI, 2023) and other text generation models still produce content that deviates from real-world facts (Bang et al., 2023; Borji, 2023; Guiven, 2023). This degrades the performance of LLMs and undermines their reliability, which is a significant bottleneck for deployment (Chuang et al., 2023; Geng et al., 2023).
Many studies have explored evaluating the factuality of LLMs Lee et al. (2022); Chuang et al. (2023); Shi et al. (2023); Chen et al. (2023). Two challenges have been identified: (i) it is difficult to assess the factuality of open-domain free-form responses, and (ii) different papers use different evaluation datasets and measures, which makes it hard to compare them, thus hampering future progress Wang et al. (2024c). To mitigate these issues, we introduce OpenFactCheck, an Open-source Factuality Evaluation Framework for LLMs.
Figure 1: Overview of the OpenFactCheck demo system for LLM factuality evaluation with three modules. Green ResponseEval: a customized fact-checker to identify factual errors given text inputs. Orange LLMEval: an LLM factuality evaluator to assess the LLM factual ability from different aspects and then to produce a report to illustrate its weaknesses and strengths. Purple CheckerEval: a fact-checker evaluator and leaderboard to encourage the development of advanced checkers in terms of performance, latency and costs.
OpenFactCheck comprises the following three core modules as shown in Figure 1:
  • ResponseEval: It allows users to customize an automatic fact-checker and to verify free-form documents to alleviate the first problem.
@@
Users can tailor their checkers according to their specific needs, such as domain specialization, cost-effectiveness, or rapid processing, and identify factual errors for both human-written text (a claim or document) and the outputs of LLMs. LLM researchers and practitioners can directly submit their LLM responses to the LLMEval by downloading our question set. Subsequently, we conduct evaluations to assess the model’s factual accuracy and to generate a report analyzing the model performance from multiple aspects. Similarly, developers who seek to evaluate and to fairly compare the efficacy of their fact-checking systems to other ones can upload their checker’s verification outcomes to CheckerEval. Then, our system will show the ranking information in the leaderboard after evaluating under the same measurements.
To sum, three modules of OpenFactCheck respectively address:
@@
We have launched an open-source initiative that includes the development of a Python library and a web interface tailored to support three major functionalities. This foundation is expected to act as a catalyst for future advancements in this domain.
We encourage extensive implementation of unique, effective, and robust claim processors, retrievers and verifiers within fact-checking pipelines, collections of challenging questions that LLMs tend to make factual errors, and human-annotated fine-grained verification examples. We believe that this will help to promote and to advance future research on LLM factuality.
@@ 2.1 ResponseEval
ResponseEval allows users to customize a fact-checking system by selecting a claim processor, a retriever, and a verifier in web pages. Current version supports the following fact-checking systems: RARR, FacTool and Factcheck-GPT (Gao et al., 2022; Chern et al., 2023; Wang et al., 2023).
@@
We consolidate various fact-checking systems into a three-step process, encapsulated by three classes: claim_processor, retriever, and verifier (Wang et al., 2024c). These classes are instantiated and sequentially connected to form a pipeline that addresses the following tasks: (i) breaking down a document into individual claims, (ii) gathering pertinent evidence for each claim, and (iii) evaluating the veracity of each claim based on the evidence provided. This sequence of tasks is referred to as solvers (see the pseudo code in Figure 4).
The implementation of a task solver can be flexible, just ensuring that the input and the output are aligned with the abstract class definitions. For example, evidence can be retrieved by calling SerpAPI or by searching Wikipedia using BM25, but we must return a list of relevant passages given an input claim. Moreover, task solvers in our pipeline are not hard-coded, but can be configured through a YAML configuration file. Thus, users can combine task-solver implementations from different systems (e.g., using Factcheck-GPT’s claim processor, RARR’s retriever, and FacTool’s verifier) and start the verification from any step. For example, users can start from the step of retrieval when the input does not need decomposition.
@@ 2.2 LLMEval
We observed that studies assessing language models’ factuality or evaluating whether the methods are effective to mitigate model hallucinations use different datasets and metrics. This makes it difficult to compare, in the same conditions, the factuality of different models as well as to compare the effectiveness of different factuality enhancement approaches. Moreover, a lot of prior work applied datasets such as MMLU (Hendrycks et al., 2021), StrategyQA (Geva et al., 2021) and HotpotQA (Yang and et al., 2018) to evaluate model’s factuality. These datasets tend to focus on assessing the general performance, rather than factuality. To this end, we first collect a dataset FactQA by gathering factual questions of existing datasets that are curated to probe diverse factual errors and span across a spectrum of domains, to fairly evaluate LLMs’ factuality under the same criteria
@@
We collected factual questions from seven commonly-used corpora that is collected deliberately to assess LLM’s factuality, including Snowball (Zhang et al., 2023a), SelfAware (Yin et al., 2023), FreshQA (Vu et al., 2023), FacTool (Chern et al., 2023), FELM-WK (Chen et al., 2023), Factcheck-GPT (Wang et al., 2023) and FactScore-Bio, a total of 6,480 examples shown in Table 1, referring to FactQA (see dataset details in Appendix B.2).
@@
Type1: Knowledge error is the most common error, occurring when the model produces hallucinated or inaccurate information due to lacking relevant knowledge or internalizing false knowledge in the pre-training stage or in the problematic alignment process.
Type2: Over-commitment error occurs when the model fails to recognize the falsehoods (or jokes) inherent in the prompt or previously-generated context, and provides an inaccurate or inappropriate response.
Type3: Disability error happens when the model is unable to search up-to-date information to correctly answer questions whose answers change over time, e.g., What is today’s gas price in New York (fast-changing). See more in Appendix B.1.
@@
For questions that can be answered by Yes/No or have a short gold answer, we perform exact matching between the model responses and the gold standard answer to judge whether the response is factually correct or not, and then to calculate accuracy, such as for Snowball and SelfAware. For FreshQA, we use the FreshEval proposed in Vu et al. (2023) to evaluate the correctness of model’s responses. For open-domain questions from the other four datasets with free-form and long responses, there are no gold standard answers. We use automatic fact-checking systems to judge the correctness of claims and obtain the percentage of true claims as the accuracy for a response.
@@ 3.1 Python Library
The first step is to import the necessary library components and initialize OpenFactCheckConfig configuration and OpenFactCheck class, which requires no input values for default usage, as shown below: {minted}python from openfactcheck.core.base import OpenFactCheck, OpenFactCheckConfig config = OpenFactCheckConfig() ofc = OpenFactCheck(config)
@@ 3.2 Web Interface
In ResponseEval page as shown in Figure 3 (b), users can click the dropdown list to select from a range of pre-implemented claim processor, retriever, and verifier. Then, users can input text either written by human or generated by machine into the text box and click Check Factuality to obtain the verification results. As the example demonstrated in the Figure, it includes two claims. The system collected 16 pieces of evidence, and one claims is supported and one claim is refuted, resulting the overall credibility of 50% and judgement “False” for this whole input.
@@
It may takes some time for LLMEval to generate teh evaluation report, depending on the system’s current load. Once the report is ready, it is emailed directly to the user, eliminating the need to wait within the application. LLM factuality evaluation report presents LLM factuality from various aspects, and specifically includes accuracy and confusion matrix of short answers, pie chart indicating accuracy over fresh questions and bar chart showing the percentage of true, false, controversial claims for free-form responses (see Figure 3 (e)).
@@ 4 Conclusion and Future Work
We implemented a unified, easy-to-use and extensible framework OpenFactCheck. It is accessible by both Python libaray and web service, supporting the customization and evaluation of automatic fact-checking systems and LLM factuality evaluation. Specifically, OpenFactCheck allows general users to check whether a claim and a document are factual or not by clicking Check, and also facilitate LLM practitioners and developers to effectively and efficiently evaluate the factuality of their LLMs from various perspectives, and to assess the accuracy of automatic fact-checking systems. In the future, we will continue to integrate new techniques, features, and evaluation benchmarks to OpenFactCheck to facilitate the research progress of LLM fact-checking.
Limitations
@@
Appendix A Related Work
While numerous automatic fact-checking systems have developed, such as RARR, FactScore, FacTool, Factcheck-GPT and Longform SAFE Gao et al. (2022); Min et al. (2023); Chern et al. (2023); Wang et al. (2023); Wei et al. (2024), they are often inaccessible to general users who lack a Python environment to compile code and run verification. Although these systems can function as the backend of a service, a user-friendly web interface is necessary to allow general users to verify text inputs by simply typing or copying text and clicking a check button. OpenFactCheck addresses this by providing an accessible web interface.
In addition, various fact-checking systems have distinct advantages. For instance, Factcheck-GPT offers a fine-grained framework to involve all possible subtask that could improve the fact-checking system, FacTool employs a low-latency evidence retriever through asynchronous processing, and FactScore introduces a scoring metric that calculates the percentage of true claims in a given text, thereby quantitatively assessing the credibility of the input. OpenFactCheck integrates these advantages into a unified system Wang et al. (2024c).
@@ Appendix A Related Work
Appendix B System Architecture
B.1 Three Common Factual Error Types
@@
Type3: Disability error happens when the model is unable to search up-to-date information to correctly answer questions whose answers change over time, e.g., What is today’s gas price in New York (fast-changing). Retrieving external knowledge and augmenting it in the context would help.
Note that we do not consider reasoning errors that arise when a claim employs flawed reasoning or faulty logic, and irrelevant error concerning that the content is unrelated to the question (Chen et al., 2023). The former highlights LLM’s reasoning ability, which is more reflected in math and reasoning tasks, and the latter has more to do with response’s helpfulness or human preference. They are important in LLM evaluation, and may implicitly influence factuality, but we will first focus on explicit causes, leaving the implicit for future work.
@@
B.2 FactQA Component Datasets

### v2

@@ Abstract
The increased use of large language models (LLMs) across a variety of real-world applications calls for automatic tools to check the factual accuracy of their outputs, as LLMs often hallucinate. This is difficult as it requires assessing the factuality of free-form open-domain responses. While there has been a lot of research on this topic, different papers use different evaluation benchmarks and measures, which makes them hard to compare and hampers future progress. To mitigate these issues, we developed OpenFactCheck, a unified framework, with three modules: (iResponseEval, which allows users to easily customize an automatic fact-checking system and to assess the factuality of all claims in an input document using that system, (iiLLMEval, which assesses the overall factuality of an LLM, and (iiiCheckerEval, a module to evaluate automatic fact-checking systems. OpenFactCheck is open-sourced [^1] https://github.com/mbzuai-nlp/openfactcheck and publicly released as a Python library [^2] https://pypi.org/project/openfactcheck/ and also as a web service [^3] http://app.openfactcheck.com. A video describing the system is available at https://youtu.be/-i9VKL0HleI.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in generating naturally-sounding answers over a broad range of human inquiries. However, GPT-4o (OpenAI, 2023) and other text generation models still produce content that deviates from real-world facts (Bang et al., 2023; Borji, 2023; Guiven, 2023). This degrades the performance of LLMs and undermines their reliability, which is a significant bottleneck for their deployment (Chuang et al., 2023; Geng et al., 2023), especially for high-stake applications, e.g., clinical, legal, and financial settings.
Figure 1: Overview of the OpenFactCheck demo system for LLM factuality evaluation and its modules. Green ResponseEval: a customized fact-checker to identify factual errors given text inputs. Orange LLMEval: an LLM factuality evaluator to assess the LLM factual ability from different aspects and then to produce a report to illustrate its weaknesses and strengths. Purple CheckerEval: a fact-checker evaluator and leaderboard to encourage the development of advanced checkers in terms of performance, latency and costs.
Many studies have explored evaluating the factuality of LLMs Lee et al. (2022); Chuang et al. (2023); Shi et al. (2023); Chen et al. (2023). Two challenges have been identified: (i) it is difficult to assess the factuality of open-domain free-form responses, and (ii) different papers use different evaluation datasets and measures, which makes it hard to compare them, thus hampering future progress Wang et al. (2024c). To mitigate these issues, we introduce OpenFactCheck.
OpenFactCheck is an Open-source Factuality Evaluation Framework for LLMs and it comprises the following three core modules (see Figure 1):
  • ResponseEval: It allows users to customize an automatic fact-checker and to verify the factuality of all claims made in a free-form document to alleviate the first problem.
@@
General users can tailor their checkers according to their specific needs, such as domain specialization, cost-effectiveness, or rapid processing, and identify factual errors for both human-written text (a claim or document) and the outputs of LLMs. LLM researchers and practitioners can directly submit their LLM responses to the LLMEval by downloading our question set. Subsequently, we conduct evaluations to assess the model’s factual accuracy and to generate a report analyzing the model performance from multiple aspects. Similarly, developers who seek to evaluate and to fairly compare the efficacy of their fact-checking systems to other ones can upload their checker’s verification outcomes to CheckerEval. Then, our system will show the ranking information in the leaderboard after evaluating under the same measurements.
To sum, three modules of OpenFactCheck respectively address the following:
@@
We have launched an open-source initiative that includes the development of a Python library and a web interface tailored to support three major functionalities. This foundation is expected to act as a catalyst for future advancements in factuality evaluation for LLMs. We encourage extensive implementation of unique, effective, and robust claim processors, retrievers and verifiers within fact-checking pipelines, collections of challenging questions that LLMs tend to make factual errors, and human-annotated fine-grained verification examples. We believe that this will help to promote and to advance future research on LLM factuality.
2 Related Work
While numerous automatic fact-checking systems have developed, such as RARR, FactScore, FacTool, Factcheck-GPT, Longform SAFE and FIRE Gao et al. (2022); Min et al. (2023); Chern et al. (2023); Wang et al. (2023); Wei et al. (2024); Xie et al. (2024), they are often inaccessible to general users who lack a Python environment to compile code and run verification. Although these systems can function as the backend of a service, a user-friendly web interface is necessary to allow general users to verify text inputs by simply typing or copying text and clicking a check button. OpenFactCheck addresses this by providing an accessible web interface.
In addition, various fact-checking systems have distinct advantages. For instance, Factcheck-GPT offers a fine-grained framework to involve all possible subtasks that could improve the fact-checking system, FacTool uses a low-latency evidence retriever through asynchronous processing, and FactScore introduces a scoring metric that calculates the percentage of true claims in a given text, thereby quantitatively assessing the credibility of the input. OpenFactCheck integrates these components into a unified system Wang et al. (2024c).
@@ 3.1 ResponseEval
ResponseEval allows users to build a customized fact-checking system by selecting a claim processor, a retriever, and a verifier in web pages. The current version supports the following fact-checking systems: RARR, FacTool and Factcheck-GPT (Gao et al., 2022; Chern et al., 2023; Wang et al., 2023).
@@
We consolidate various fact-checking systems into a three-step process, encapsulated by three classes: claim_processor, retriever, and verifier (Wang et al., 2024c). These classes are instantiated and sequentially connected to form a pipeline that addresses the following tasks: (i) breaking down a document into individual claims, (ii) gathering pertinent evidence for each claim, and (iii) evaluating the veracity of each claim based on the evidence provided. This sequence of tasks is referred to as solvers. (see the pseudo code in Appendix A)
The implementation of a task solver can be flexible, just ensuring that the input and the output are aligned with the abstract class definitions. For example, evidence can be retrieved by calling SerpAPI or by searching Wikipedia using BM25, but we must return a list of relevant passages given an input claim.
Moreover, task solvers in our pipeline are not hard-coded, but can be configured through a YAML configuration file. Thus, users can combine task-solver implementations from different systems (e.g., using Factcheck-GPT’s claim processor, RARR’s retriever, and FacTool’s verifier) and start the verification from any step. For example, users can start from the step of retrieval when the input does not need decomposition.
@@ 3.2 LLMEval
We observed that studies assessing language models’ factuality or evaluating whether the methods are effective to mitigate model hallucinations use different datasets and metrics. This makes it difficult to compare, in the same conditions, the factuality of different models as well as to compare the effectiveness of different factuality enhancement approaches. Moreover, a lot of prior work applied datasets such as MMLU (Hendrycks et al., 2021), StrategyQA (Geva et al., 2021) and HotpotQA (Yang et al., 2018) to evaluate model’s factuality. These datasets tend to focus on assessing the general performance, rather than factuality. To this end, we first collect a dataset FactQA by gathering factual questions of existing datasets that are curated to probe diverse factual errors and span across a spectrum of domains, to fairly evaluate LLMs’ factuality under the same criteria
@@
We collected factual questions from seven commonly-used corpora that is collected deliberately to assess LLM’s factuality, including Snowball (Zhang et al., 2023a), SelfAware (Yin et al., 2023), FreshQA (Vu et al., 2023), FacTool (Chern et al., 2023), FELM-WK (Chen et al., 2023), Factcheck-GPT (Wang et al., 2023) and FactScore-Bio, a total of 6,480 examples shown in Table 1, referring to FactQA (see dataset details in Appendix C).
@@
Type1: Knowledge error is the most common error when the model produces hallucinated or inaccurate information due to lacking relevant knowledge or internalizing false knowledge in the pre-training stage or in the alignment process.
Type2: Over-commitment error occurs when the model fails to recognize the falsehoods (or jokes) in the prompt or previously-generated context, and provides an inaccurate or inappropriate response.
Type3: Disability error happens when the model is unable to search up-to-date information to correctly answer questions whose answers change over time, e.g., What is today’s gas price in New York (fast-changing). See more in Appendix B.
@@
For questions that can be answered by Yes/No or have a short gold answer, we perform exact matching between the model responses and the gold standard answer to judge whether the response is factually correct, and then to calculate accuracy, such as for Snowball and SelfAware.
For FreshQA, we use the FreshEval proposed in Vu et al. (2023) to evaluate the correctness of model’s responses. For open-domain questions from the other four datasets with free-form and long responses, there are no gold standard answers. We use automatic fact-checking systems to judge the correctness of claims and obtain the percentage of true claims as the accuracy for a response.
@@ 3.3 CheckerEval
Discussion about Unifying
It can be argued that the underlying philosophies of the three modules differ, reflecting varying interpretations of factuality. For example, the design view o LLMEval and CheckerEval differs from that of ResponseEval.
Our goal is to integrate all LLM factuality evaluation functionality into a unified framework, while preserving the individual function.
The LLMEval employs different metrics across datasets. This may be debated. Similarly, we aim to consolidate these datasets into a unified benchmark, enabling other studies to utilize a standardized evaluation function. This approach would enhance the fairness of comparisons across studies, as they would be evaluated consistently, despite the use of dataset-specific evaluation measures.
We acknowledge the limitation of the current CheckerEval, which is restricted to evaluating the verification step. We plan to progressively extend its capabilities to support fine-grained evaluations across multiple steps. [^4] The evaluator currently supports both claim-level and document-level verification, depending on whether users download claim or document datasets.
@@ 4.1 Python Library
The first step is to import the necessary library components and initialize OpenFactCheckConfig configuration and OpenFactCheck class, which requires no input values for default usage, as shown below: {minted}python from openfactcheck import OpenFactCheck, OpenFactCheckConfig config = OpenFactCheckConfig() ofc = OpenFactCheck(config)
@@
User is provided with the benchmarks for the LLM and FactChecker evaluations, and can upload the responses to the library for evaluation in the form of CSV files. CSV file format for LLM evaluation has two columns: index and response, while the FactChecker evaluation CSV file format has three columns: label, time, and cost.
@@ 4.2 Web Interface
In ResponseEval page as shown in Figure 3 (b), users can click the dropdown list to select from a range of pre-implemented claim processor, retriever, and verifier. Then, users can input text either written by human or generated by machine into the text box and click Check Factuality to obtain the verification results. As the example demonstrated in the Figure, it includes two claims.
The system collected sixteen pieces of evidence, and one claims is supported and one claim is refuted, resulting the overall credibility of 50% and judgement “False” for this whole input.
@@
It may takes some time for LLMEval to generate teh evaluation report, depending on the system’s current load. Once the report is ready, it is emailed directly to the user, eliminating the need to wait within the application. LLM factuality evaluation report presents LLM factuality from various aspects, and specifically includes accuracy and confusion matrix of short answers, pie chart indicating accuracy over fresh questions and bar chart showing the percentage of true, false, controversial claims for free-form responses, as shown on Figure 3 (e).
@@ 5 Conclusion and Future Work
We implemented a unified, easy-to-use and extensible framework OpenFactCheck. It is accessible by both Python library and web service, supporting the customization and evaluation of automatic fact-checking systems and LLM factuality evaluation. Specifically, OpenFactCheck allows general users to check whether a claim and a document are factual or not by clicking Check, and also facilitate LLM practitioners and developers to effectively and efficiently evaluate the factuality of their LLMs from various perspectives, and to assess the accuracy of automatic fact-checking systems. In the future, we will continue to integrate new techniques, features, and evaluation benchmarks to OpenFactCheck to facilitate the research progress of LLM fact-checking.
Limitations and Future Work
@@
Multilingual Expansion
OpenFactCheck is a platform that combines the features of various fact-checking systems and is designed to be language-agnostic. While the default task solvers in the system are configured for English, the platform can be expanded to accommodate other languages by developing task solvers that align with the specific linguistic requirements of those languages. This flexibility allows for easy adaptation and extension to support multilingual fact-checking capabilities.
@@
Temporal Issues
The factuality of statements can change over time due to evolving events, new discoveries, or updated information. OpenFactCheck does not explicitly account for temporal dynamics as of now, which may lead to discrepancies between the evaluation results and the current state of knowledge. Authors are already working on factuality evaluation methods that consider temporal aspects, which will be integrated into OpenFactCheck in future releases.
@@
Appendix A Pseudo Code of ResponseEval
In this section, we present the pseudo code for the ResponseEval, a modular system designed to process, retrieve, and verify claims found in textual documents. The system is divided into three primary components: the claim processor, the retriever, and the verifier. Each module is tasked with a specific function—extracting claims from the input document, retrieving relevant evidence, and verifying the factual accuracy of the claims, respectively. Figure 4 outlines the pseudo code implementation of each module, showcasing the flow from document processing to final verification. This structured approach allows for a systematic handling of claims, leveraging both natural language processing tools and deep learning models to ensure a comprehensive evaluation of document veracity.
@@ Appendix A Pseudo Code of ResponseEval
Appendix B Factual Error Evaluation
@@
Type3: Disability error happens when the model is unable to search up-to-date information to correctly answer questions whose answers change over time, e.g., What is today’s gas price in New York (fast-changing). Retrieving external knowledge and augmenting it in the context would help for such cases. Note that we do not consider reasoning errors that arise when a claim is based on flawed reasoning or faulty logic.
Thus, ex exclude irrelevant error concerning that the content is unrelated to the question (Chen et al., 2023). The former highlights LLM’s reasoning ability, which is more reflected in math and reasoning tasks, and the latter has more to do with response’s helpfulness or human preference. They are important in LLM evaluation, and may implicitly influence factuality, but we will first focus on explicit causes, leaving the implicit for future work.
@@
Appendix C FactQA Component Datasets